OpenStudy (anonymous):

Please Help!!!!!! Solve the equations for the solutions, x and y . 2x + y = 3 x - 2y = -1 The lines intersect at __, __

OpenStudy (anonymous):

Let's use the substitution method. We'll solve for x in the second equation, then substitute it back in the first equation to solve for y, finally substituting that y value in for x. x -2y = -1 x -2y +2y = -1 +2y x = 2y -1 Great, now we have x. Let's put it into the first equation. 2x +y = 3 2(2y -1) +y = 3 4y -2 +y = 3 5y -2 = 3 5y -2 +2 = 3 +2 5y = 5 5y/5 = 5/5 y = 1 There's your y value. Now, we already know that x = 2y -1. Let's plug y in there to get x. x = 2y -1 x = 2(1) -1 x = 2 -1 x = 1 There's x. Now, plug y = 1 and x = 1 into a coordinate pair, and you have your intersection point.